OpenStudy (anonymous):

]Choose the equation of the vertical line passing through the point (-4, 2). y = -4 x = -4 y = 2 x = 2

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of . thanks ure good how bout this one

OpenStudy (anonymous):

For vertical line, we know that it must be parallel to y-axis and thus it passes through x-axis so we can find that the answer should be x=...... as it passes through x-axis For a line passing though (-4, 2), Reply Using Drawing we can know that the answer is x=-4

OpenStudy (anonymous):



OpenStudy (anonymous):

y = -4 x = -4 y = 2 x = 2

OpenStudy (anonymous):

answer is oone of them though

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of WHat is the slope given

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

y-(-4) / x-1 =1/2 y+4 / x-1 =1/2 2(y+4) = (x-1) 2y+8=x-1 x-2y-9=0

OpenStudy (anonymous):

wat?? thats not a choice

OpenStudy (anonymous):

What are the choices?